# Neural Architecture Search via Two Constant Shared Weights Initialisations

### v2

@@ 3.4 Embedding initialisation
In order to verify the effect of the embedding initialisation, we have run tests with $6$ different initialisations:
@@
Care should be taken when initialising the networks containing the embedding layer: if embedding is initialised with all constants, there is no difference between the embedded input. The performance of our metric, in this case, will be analogous to that with a batch size of one.
Our ablation studies show that as soon as the embedding is initialised with non-constant weights, it does not significantly influence the outcomes.
@@ 3.5 Synthetic data
  • random normal: images are filled with random values following normal distribution with $[\mu,\sigma]=[0,1]$
@@
Curiously, greyscale images, filled with constant values, show the closest results to the actual data. Note that epsilon metric with greyscale data outperforms synflow. It is an essential achievement because synflow does not use input data and is, therefore, data independent. Our results show that epsilon has the potential to be used with no data whatsoever.
@@
4 Visualisation of other zero-cost NAS metrics
4.1 When integrated with other NAS methods
@@
In Section 2.2, we show how epsilon metric improves the performance of ageing evolution and random search when used for warming up. Figure 5 compares epsilon integration to other metrics from Abdelfattah et al. (2021). For both EA ans RS, we use the metrics for a warm-up and run the procedure until the number of trained architectures reaches $300$ with $100$ random rounds. The warm-up pool contains $3000$ randomly selected architectures.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
@@
4.2 Correlation with accuracy
In Abdelfattah et al. (2021) the metrics are presented through statistical measures, but we feel that visualisation helps to improve understanding. Here we provide visualisations for two search spaces built on the data provided by the authors.
@@ 4.2 Correlation with accuracy
Figure 6: Zero-cost metrics performance illustration for NAS-Bench-201 search space evaluated on CIFAR-10, CIFAR-100 and ImageNet16-120 datasets, based on data from Abdelfattah et al. (2021). Each dot corresponds to an architecture; the darker the colour, the more parameters it contains. Figure represents search space of $15{,}625$ networks (excluding architectures with NaN scores).
5 Discussion
The epsilon metric performs surprisingly well on various benchmark datasets. Similarly to other NAS algorithms, it shows the most robust performance with the NAS-Bench-201 benchmark. This fact indicates that the set of architectures comprising NAS-Bench-201 forms a relatively easy-to-navigate landscape regardless of the metric. Still, epsilon shows the best performance among trainless metrics. Table 1 shows that synflow shows comparable performance in terms of correlations on the ImageNet-16-120 dataset. However, statistically epsilon selects, on average, a better-performing architecture, which can also be confirmed visually by comparing Figures 1 and 6.
Notably, we would like to highlight the difference with our previous work (cv metric, Gracheva (2021)). The two methods differ in how the weights are de-emphasised and what outputs we use for statistics computation. Two constant shared weight initialisations better bring out the effect of the bare geometry compared to a hundred random initialisations. Moreover, epsilon metric takes significantly less time to compute, being fifty-fold faster.
It is important to note that we used NAS-Bench-201 for the metric development; therefore, the results of this part of the study might be biased.
To confirm the generalisability of the metric, we applied it to the NAS-Bench-101 benchmark. Again, Table 3 shows strong performance. The peculiar vertical lines are due to clustering in terms of test accuracy. Since every architecture in NAS-Bench-101 is trained with $3$ random seeds, the four clusters correspond to situations when the training diverges for some random initialisations. More random seeds would reveal a smoother picture and better correlations. Again, we observe comparable performance between synflow and epsilon scoring metrics. We could only confirm some of the results from Abdelfattah et al. (2021) and suppose that some statistics might have been miscalculated.
By design, we expect our metric to be independent of the specific neural architecture structure. To confirm this claim, this is essential to show its applicability to architectures other than CNNs. Our work tests epsilon applicability to RNN models with NAS-Bench-NLP.
While statistics in Table 4 are not exceptional, Figure 2 shows a visible trend towards better architectures with increasing metric value. Significant noise levels might stem from the characteristics of the benchmark itself: factors like a relatively small sample of networks given a vast space, choice of training hyperparameters, dropout rates, and others may distort NAS performance. Nonetheless, the trend is strong enough to conclude that epsilon metric can apply to recurrent type architectures.
While epsilon metric shows solid empirical performance, the underlying reasons for this are unclear.
There are several hints towards its understanding. First, mathematically, epsilon represents the difference in the output distribution shapes between initialisations. The output shape is affected by layer widths, activation functions, batch normalisation, skip connections and other factors, which we generally call network geometry. By setting weights to a constant shared value, one can probe the effects of the geometry without being obstructed by the randomness of initialisation.
Second, according to the weight ablation studies (Section 3), the best performance is achieved when we set the weights to the lowest and highest values that do not lead to excessive outputs explosion or vanishing. Therefore, epsilon measures the amplitude of the outputs’ distribution shape change due to geometry.
6 Conclusion
This work presents a novel zero-cost NAS scoring metric epsilon. It allows evaluating a neural architecture’s potential based on the outputs’ statistics after two constant shared weights initialisations. The value of epsilon reflects how neural network outputs’ distribution changes between low and high constant shared weight initialisations. It shows that the higher this difference, the better the network will perform upon training.
It takes a fraction of a second to evaluate a single architecture and does not require gradients computation (low demand on memory) or labels, and can run on a CPU. We provide experimental evidence that the introduced epsilon metric generalises to various architecture types, such as convolutional or recurrent neural networks (care should be taken with embedding layer initialisation).
This work evaluates epsilon on three staple NAS search spaces: NAS-Bench-201, NAS-Bench-101 and NAS-Bench-NLP. It sets the state-of-the-art among zero-cost NAS metrics and can be easily implemented within other methods, such as random search or evolutionary algorithms.
The only significant disadvantage of the method is that it requires the choice of constant weight values used for initialisation. Our tests show that it must be set up individually for each search space. We plan to automate the weight selection process in our future work.
Overall, while a coherent theoretical foundation of epsilon is missing, we consider the empirical results compelling.

### v3

@@ 6.4 Embedding initialisation
To verify the effect of the embedding initialisation, we have run tests with $6$ different initialisations:
@@
Care should be taken when initialising the networks containing the embedding layer: if embedding is initialised with all constants, there is no difference between the embedded input. In this case, our metric’s performance will be analogous to that with a batch size of one.
Our ablation studies show that it does not significantly influence the outcomes as soon as the embedding is initialised with non-constant weights.
@@ 6.5 Synthetic data
  • random normal: images are filled with random values following a normal distribution with $[\mu,\sigma]=[0,1]$
@@
Curiously, greyscale images, filled with constant values, show the closest results to the actual data. Note that epsinas metric with greyscale data outperforms synflow. It is an essential achievement because synflow does not use input data and is, therefore, data independent. Our results show that epsinas has the potential to be used with no data whatsoever.
@@
7 Discussion
Our experiments show that despite its simplicity, epsinas metric shows surprisingly good performance across different benchmark datasets. It shows the most consistent performance with NAS-Bench-201, suggesting that this benchmark defines a relatively easy-to-navigate search space. Tables 2, 4, 5 demonstrates that epsinas outperforms most of other zero-cost metrics in terms of Spearman and Kendal correlations. The differences are statistically significant, with p-values between epsinas and other metrics below 0.001.
Notably, we highlight the key differences between our previous work (cv metric, [18]) and the current approach. The two methods differ in how they de-emphasize weights and the type of outputs used for statistical computation. Using two fixed weight initialisations more effectively isolates the influence of network geometry compared to averaging over a hundred random initialisations. Additionally, the epsinas metric is significantly more efficient, reducing computation time by a factor of fifty.
We also compare the performance of epsinas with its conceptual counterparts, zico and $\xi$-gsnr. While the metrics are based on similar principles, zico and $\xi$-gsnr use gradients instead of raw outputs, which appears to limit their performance. This difference may also be attributed to random initialisation, as opposed to the shared constant weight initialisation used by epsinas. In all cases, epsinas consistently outperforms the naive nparams proxy, which relies solely on the total number of parameters in a given network.
It is important to note that we used NAS-Bench-201 to develop our metric, so the findings from this part of the study may be biased. To test the generalizability of the metric, we applied it to the NAS-Bench-101 benchmark. As shown in Table 4, epsinas continues to demonstrate strong performance. The peculiar vertical lines in the results are due to clustering based on test accuracy. Since each architecture in NAS-Bench-101 is trained with three random seeds, the four clusters represent cases where training diverges for certain random initialisations. Using more random seeds would likely reveal a smoother pattern and stronger correlations.
By design, we expect our metric to be independent of the specific neural architecture structure. To confirm this claim, we show its applicability to architectures to RNN models with NAS-Bench-NLP.
While the statistics in Table 5 may not seem extraordinary at first glance, Figure 2 demonstrates a visible trend of improved architecture selection as the epsinas metric value increases. This positive trend is particularly noteworthy, especially considering the noise inherent to the benchmark. Factors such as the limited sample size of networks within a vast architecture space, variations in training hyperparameters, dropout rates, and other variables can contribute to performance inconsistencies in NAS. Despite these challenges, the trend is strong and consistent enough to confidently conclude that the epsinas metric shows great potential for effectively evaluating recurrent-type architectures.
While epsinas metric shows solid empirical performance, the underlying reasons for this remain unclear.
There are several hints towards its understanding. First, mathematically, epsinas represents the difference in the output distribution shapes between initialisations. The output shape is affected by layer widths, activation functions, batch normalisation, skip connections and other factors, which we generally call network geometry. By setting weights to a constant shared value, one can probe the effects of the geometry without being obstructed by the randomness of initialisation.
Second, according to the weight ablation studies (Section 6), the best performance is achieved when we set the weights to the lowest and highest values that do not lead to excessive output explosion or vanishing. Therefore, epsinas measures the amplitude of the outputs’ distribution shape change due to geometry.
7.1 Future directions
While the presented epsinas metric demonstrates solid performance, several open questions remain. One of the key directions for future development is to assess its applicability to more types of neural topologies, such as transformer [3, 54], generative adversarial [17] or graph neural networks [39], and to test it with hierarchical search spaces [6, 39, 29, 3]. It is also interesting to evaluate the metric in combination with other advanced NAS algorithms, such as multi-objective optimization [53], differentiable architecture search [50], Bayesian optimization [43, 42], or within the Parametric Zero-Cost Proxies [10] framework.
The metric’s performance is influenced by various hyperparameters, including weight values, batch size, and initialization scheme. While we conducted ablation studies on their individual effects, potential interactions between these hyperparameters remain unexplored, making it important to investigate their joint impact. Additionally, the current weight setup is based on empirical choices and would benefit from a solid theoretical foundation.
A deeper theoretical study of the metric is particularly needed given its similarities to the zico metric. For instance, it remains unclear why the mean absolute error over raw outputs outperforms the deviation over gradients and how this behaviour relates to the initialization scheme. Furthermore, examining the performance of the zico metric when initialized with constant shared weights could provide further insights.
8 Conclusion
This work presents a simple zero-cost NAS scoring metric, epsinas, which measures how much a network’s outputs change when initialized with two different constant shared weights. We compute it by performing two forward passes and calculating the normalized MAE between the raw outputs.
For a simple metric, epsinas shows an unexpectedly high correlation with network performance in CNNs and RNNs, outperforming most existing zero-shot NAS proxies. We evaluate the epsinas metric on three widely-used NAS search spaces: NAS-Bench-201, NAS-Bench-101, and NAS-Bench-NLP. Our results show strong performance across these search spaces.
The metric is computationally efficient, requiring only a fraction of a GPU second to evaluate a single architecture, with minimal memory demand, as it does not require gradient computation or labelled data. epsinas can be run on a CPU, making it broadly accessible. While small to medium-sized networks can be processed efficiently on a CPU, larger architectures may require GPU acceleration to handle forward passes more quickly, especially when evaluating multiple candidate architectures in succession. Optimizing CPU-GPU workload distribution could help balance energy consumption and speed in resource-constrained environments such as mobile and embedded systems. In our implementation, we have already incorporated a degree of heterogeneous computing by offloading output data from the GPU to the CPU for statistical computation. Such optimisations could be valuable in domains where rapid evaluation of complex architectures is necessary.
Experimental results demonstrate that epsinas applies to convolutional networks and generalizes well to recurrent networks—with careful consideration for embedding layer initialization. The method can also be readily incorporated into other NAS frameworks, such as random search and evolutionary algorithms.
The main downside of our metric is that the optimal weight values need to be set separately for different search spaces and neural topologies. We found that using both high and low weight values helps highlight differences in initialization, leading to better correlations with final performance. While the weights should cover both extremes to test the architecture’s robustness, it’s unclear how to automate this selection process.
These results offer an intriguing clue about the inner workings of neural networks. We hope that epsinas ’s performance can be improved with a better theoretical understanding of the optimal weight setup. We believe that further development of this method will lead to a faster and computationally efficient NAS and a clearer understanding of network prediction and generalization mechanisms.
These results might offer some interesting clues about how neural networks work. We are hopeful that with a better theoretical understanding of the optimal weight setup, epsinas ’s performance can improve, and we believe that further development of this method could help us better understand network prediction and generalization.
Declarations
Competing interests
The authors have no competing interests to declare that are relevant to the content of this article.
Funding
No funding was received for conducting this study.
Code availability
The code supporting this study can be found on GitHub at https://github.com/egracheva/epsinas.
Appendix A Visualisation of other zero-cost NAS metrics
A.1 When integrated with other NAS methods
@@
In Section 5.6, we show how epsinas metric improves the performance of ageing evolution and random search when used for warming up. Figure 7 compares epsinas integration to other metrics from [1]. For both EA and RS, we use the metrics for a warm-up and run the procedure until the number of trained architectures reaches $300$ with $100$ random rounds. The warm-up pool contains $3000$ randomly selected architectures.
@@
A.2 Correlation with accuracy
In [1], the metrics are presented through statistical measures, but we feel that visualisation helps to improve understanding. Here, we provide visualisations for two search spaces built on the data provided by the authors.
@@ A.2 Correlation with accuracy
Figure 8: Zero-cost metrics performance illustration for NAS-Bench-201 search space evaluated on CIFAR-10, CIFAR-100 and ImageNet16-120 datasets, based on data from [1]. Each dot corresponds to an architecture; the darker the colour, the more parameters it contains. The figure represents the search space of $15{,}625$ networks (excluding architectures with NaN scores).